$740 is split into two investments, one paying 6% and the other paying %9 interest. if the total amount of money earned from the two investments is $52.50, how much money is invested at each rate

Answer by macston(5194)

S=invested at 6%; N=invested at 9%
S+N=$740
S=$740-N
0.06S+0.09N=$52.50 Substitute for S
0.06($740-N)+0.09N=$52.50
$44.40-0.06N+0.09N=$52.50 Subtract $44.40 from each side.
0.03N=$8.10 Divide each side by 0.03
N=$270 ANSWER 1: The amount invested at 9% was $270.
S=$740-$270=$470 ANSWER 2: The amount invested at 6% was $470.
CHECK
0.06S+0.09N=$52.50
0.06($470)+0.09($270)=$52.50
$28.20+$24.30=$52.50
$52.50=$52.50